OpenStudy (anonymous):

Find the partial fraction decomposition for the rational expression. 15x-21 OVER x^2-3x+2

OpenStudy (anonymous):

x^2-3x+2 can be factored

OpenStudy (anonymous):

? what?

OpenStudy (anonymous):

Factor \(x^2 - 3x + 2\) into \((x-2)(x-1)\)

OpenStudy (anonymous):

o how u knoow how to do that

OpenStudy (anonymous):

Practice

OpenStudy (anonymous):

how did u do that ha

OpenStudy (anonymous):

\[(x + a)(x + b) = x^2 + (a+b)x + ab\] Therefore you are looking for an a and b such that: \[a+b = -3 \text{ and } a\cdot b = 2\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

still not done yes

OpenStudy (anonymous):

\[\frac{15x-21}{(x-2)(x-1)}=\frac{A}{x-2}+\frac{B}{x-1}\]

OpenStudy (anonymous):

\[A=\frac{15\times 2-21}{(2-1)}=9\]

OpenStudy (anonymous):

\[B=\frac{15\times 1-21}{1-2}=6\]

OpenStudy (anonymous):

answer is \[\frac{15x-21}{(x-2)(x-1)}=\frac{9}{x-2}+\frac{6}{x-1}\]